OpenStudy (amtran_bus):

if the radius of the CD is 6.0 cm and the angular velocity is 5.0 rad/s, what is the linear displacement of a point on the edge in 2.0 s?

OpenStudy (amtran_bus):

@guillefix @ghazi

OpenStudy (ghazi):

well v=r*w v- linear velocity v= 0.06*5=0.3 m/s so linear displacement is d=velocity * time = 0.3*2=0.6 m

OpenStudy (amtran_bus):

How did you get the 0.3 m/s? @ghazi

OpenStudy (amtran_bus):

Ok, I see, thanks.

OpenStudy (amtran_bus):

Could you also help me find the angular displacement?

OpenStudy (ghazi):

well angular displacement is theta = w* time = 5*2= 10 rad

OpenStudy (amtran_bus):

Your an awesome help.

OpenStudy (ghazi):

you are welcome :D